# Simplifying Debiased Inference via Automatic Differentiation and Probabilistic Programming

### v1

@@ Abstract
Keywords: efficient influence function, semiparametric efficiency, pathwise differentiability, Hadamard differentiability, automatic differentiation, probabilistic programming, differentiable programming
@@ 1.1 Motivation
There is a long history of constructing estimators that are asymptotically efficient (‘efficient’ hereafter) in infinite-dimensional models (Levit, 1975; Hasminskii and Ibragimov, 1979; Pfanzagl, 1990; Bickel et al., 1993). Several approaches for doing this permit the use of arbitrary methods for estimating needed nuisance functions. These include one-step estimation (Pfanzagl, 1982), estimating equations (van der Laan et al., 2003), targeted learning (van der Laan and Rubin, 2006), and double machine learning (Chernozhukov et al., 2017). These frameworks construct efficient estimators using an object known as the efficient influence function (EIF). Though EIFs have been derived in many examples (Newey, 1990; Bickel et al., 1993; van der Laan et al., 2003), computing them in new problems can be challenging, requiring tools from functional analysis and differential geometry. These calculations can represent a bottleneck in scientific discovery, as some novel analyses cannot be performed until an expert derives the EIF.
@@
Automatic differentiation avoids these challenges (Wengert, 1964; Speelpenning, 1980). In this approach, a function $f$ is written as a composition of differentiable functions with known gradients, which we refer to as primitives. Then, an algorithm can iteratively apply the chain rule to derive the gradient of $f$ at a specified point. Unlike numerical differentiation, automatic differentiation is hyperparameter-free and is always accurate up to working precision. Differentiable programming extends automatic differentiation by identifying a compositional structure for $f$ directly via computer code that can be used to evaluate it, simplifying implementation (Baydin et al., 2018; Blondel and Roulet, 2024). Automatic differentiation and differentiable programming have been used with success in applications (Bischof, 1995; Homescu, 2011), and have played a crucial role in the rapid progress made in deep learning (Goodfellow et al., 2016). However, they have yet to be used to compute an EIF. This may be due to the fact that EIFs correspond to a gradient of a functional defined on an infinite-dimensional statistical manifold, and so it has been unclear what primitives can be chained together to automate their computation.
@@ 1.2 Contributions
    We provide an automatic differentiation algorithm for computing the EIF of any parameter that can be expressed as a composition of Hilbert-valued primitives.
@@
    We compile a table of primitives that can express many parameters of interest.
We then turn to estimation, building on developments in probabilistic programming (van de Meent et al., 2018). Probabilistic programming has been successfully employed in Bayesian inference (Ge et al., 2018; Tran et al., 2018; Stan Development Team, 2023), parametric maximum likelihood estimation (Fournier et al., 2012; Kristensen et al., 2015) and M-estimation (Zivich et al., 2022; Kosmidis and Lunardon, 2024). However, to date there is not a probabilistic programming framework for efficiently estimating a generic smooth parameter that may rely on nuisances like density and regression functions. Our remaining contributions address this gap.
@@
    We introduce an algorithm to estimate any parameter that is expressed as a composition of known primitives and conditions under which this estimator will be efficient.
@@
    We provide a proof-of-concept Python package implementing our approach, available at http://github.com/alexluedtke12/dimple.
@@
The goal of dimple is similar in spirit to that of automatic debiased machine learning methods: to avoid analytical derivations when constructing efficient estimators (Chernozhukov et al., 2022a, b, c). Those existing methods apply to regression functionals, whose EIFs depend on the functional’s form through nuisance functions known as Riesz representers. Automatic debiased machine learning provides loss functions that can be used to estimate these nuisances without explicitly deriving their form. Compared to these methods, dimple innovates in several directions. It allows the functional to depend on the distribution through non-regression summaries, uses automatic differentiation to compute the EIF, and is integrated into a user-friendly probabilistic programming framework. It also works in semiparametric as well as nonparametric models and applies to general Hilbert-valued parameters (Luedtke and Chung, 2023).
Beyond simplifying estimation, our automatic differentiation approach and the chain rule that underlies it may be of independent interest for their use in deriving semiparametric efficiency bounds, which are given by the variance of the EIF when it exists (van der Vaart and Wellner, 1989; van der Vaart, 1991a). Our theoretical results show this EIF exists for any parameter expressible as a composition of differentiable primitives, and applying automatic differentiation by hand gives its analytic form. See Section 2.3 for an example.
@@ 1.3 Illustrations of proposed approach
Our proof-of-concept Python package uses dimple to construct nonparametric efficient estimators. We illustrate it with three examples. The following Python code imports the classes and functions from the package needed to run them:
from dimple import Distribution, E, Density, Var, RV, estimate
@@ 1 P = Distribution(data=dat)
The first line of code defines a distribution $P$ and identifies a dataset dat containing iid draws from it to be used for estimation. The next two lines serve as a blueprint of the statistical estimation problem, formulating it using primitives from the dimple package. The fourth line computes a 5-fold cross-fitted estimator, Wald-type standard error, and 95% confidence interval of $\psi(P)$. Internally, the dimple package estimates density functions using kernel density estimation and all other nuisances using lightgbm (Ke et al., 2017), with hyperparameters tuned via Optuna (Akiba et al., 2019).
@@ 4 R2 = 1-E(P,(RV(’Y’)-mu)**2)/v # $\psi(P)$
The final example concerns the estimation of the longitudinal G-formula, which corresponds to the mean outcome under a longitudinal binary treatment. This parameter is defined in longitudinal settings where covariate-treatment pairs $(X_{t},A_{t})$ are collected at times $t=0,1,\ldots,T-1$, and then an outcome $Y$ is collected at the end of the study. For $t\in\{0,1,\ldots,T-1\}$, let $H_{t}$ denote the history $(X_{0},A_{0},X_{1},A_{1},\ldots,X_{t-1},A_{t-1},X_{t})$ available just before treatment $A_{t}$ at time $t$ is administered. Also let $H_{T}=(H_{T-1},A_{T},Y)$ denote all of the covariate, treatment, and outcome data. The parameter $\psi(P)$ is defined recursively by letting $\mu_{P,T}(h_{T})=y$, $\mu_{P,t}(h_{t})=E_{P}[\mu_{P,t+1}(H_{t+1})\,|\,A_{t}=1,H_{t}=h_{t}]$ for $t=T-1,T-2,\ldots,0$, and $\psi(P)=E_{P}[\mu_{P,0}(H_{0})]$. The recursive definition of this parameter can be easily expressed in dimple via a for loop, as illustrated below.
@@ 5 H = [f’X{j}’ for j in range(t+1)]+[f’A{j}’ for j in range(t)] # $H_{t}$
Example 3: Longitudinal G-Formula (Robins, 1986; Bang and Robins, 2005; Luedtke et al., 2017; Rotnitzky et al., 2017; Chernozhukov et al., 2022b).
In Appendix A, we evaluate the performance of dimple in a Monte Carlo study. The results support that dimple can produce asymptotically efficient estimators and valid confidence intervals. At sample sizes of $n\geq 1000$, the confidence intervals have approximately nominal coverage and nearly optimal widths, scaling with the standard deviation of the EIF. These results are not surprising given that dimple uses a variant of one-step estimation, which has been shown to perform well in a variety of settings (Pfanzagl, 1982; Bickel et al., 1993). Details on this estimation framework and its theoretical properties are given in Section 3.1.
@@ 2.1 Algorithm and theoretical guarantee
Though we have not seen this precise notion of differentiability introduced previously, total pathwise differentiability is closely related to classical notions that apply to functions of a single argument, $u\in\mathcal{U}$ or $P\in\mathcal{M}$. In particular, the total pathwise differentiability of $\theta$ at $(P,u)$ implies $\theta(P,\,\cdot\,)$ is Hadamard differentiable (Averbukh and Smolyanov, 1967) and $\theta(\,\cdot\,,u)$ is pathwise differentiable (van der Vaart, 1991a), and the converse implication holds under additional regularity conditions — see Lemmas S5 and S6 in the appendix. In the special case where $\mathcal{U}$ is the trivial vector space $\{0\}$, $\theta$ is totally pathwise differentiable if and only if $\nu(\cdot):=\theta(\,\cdot\,,0)$ is pathwise differentiable. In these cases, $\dot{\theta}_{P}^{*}(w)=(\dot{\nu}_{P}^{*}(w),0)$. The map $\dot{\nu}_{P}^{*}:\mathcal{W}\rightarrow\dot{\mathcal{M}}_{P}$ is called the efficient influence operator of $\nu$ at $P$; it can be used to draw inferences about $\nu(P)$ (Luedtke and Chung, 2023). When $\mathcal{W}=\mathbb{R}$ and $w=1$, $\dot{\nu}_{P}^{*}(w)$ is known as the EIF of $\nu$ at $P$.
Like Hadamard and pathwise differentiability (Averbukh and Smolyanov, 1967; van der Vaart, 1991b), total pathwise differentiability satisfies a chain rule — see Lemma S3 in the appendix. The availability of a chain rule suggests a form of automatic differentiation (Linnainmaa, 1970), which we present in Algorithm 2. Initially, this algorithm requires the evaluation of the composition that defines $\psi$. Subsequently, information backpropagates through the indices $j$ of the composition to evaluate the efficient influence operator. This information is derived from the adjoint $\dot{\theta}_{j,P,h_{\mathrm{pa}(j)}}^{*}$ of the differential operator of $\theta_{j}$ at $(P,h_{\mathrm{pa}(j)})$, where $h_{\mathrm{pa}(j)}$ is as defined as in Algorithm 1 when that algorithm is called with input $P$.
@@ 2.2 Useful primitives
Table 1: Primitives $\theta$ with domain ×$\mathcal{M}\times\mathcal{U}$. Given the specified ambient Hilbert spaces $\mathcal{T}\supseteq\mathcal{U}$ and $\mathcal{W}\supseteq\mathrm{Image}(\theta)$, each is totally pathwise differentiable under conditions given in Appendix D.
@@
Table 1 provides a selection of useful primitives. In Appendix D, we provide sufficient conditions under which each is totally pathwise differentiable when $\mathcal{M}$ is a locally nonparametric model of distributions on $\mathcal{Z}$. Some of these primitives involve a coarsened random variable $X=\mathscr{C}(Z)$, where $\mathscr{C}:\mathcal{Z}\rightarrow\mathcal{X}$ is a many-to-one map; when $Z\sim Q\in\mathcal{M}$, we let $Q_{X}$ denote the marginal distribution of $X$. The primitives fall into three distinct categories, which we elaborate on below.
@@
The choice of ambient Hilbert spaces $\mathcal{T}$ and $\mathcal{W}$ can impact a primitive’s total pathwise differentiability. For the first five primitives in Table 1 and the conditional density primitive, this choice involves selecting a $Q\in\mathcal{M}$ indexing one or more $L^{2}$ spaces. As detailed in Appendix D, sufficient conditions for each primitive’s total pathwise differentiability at a given $(P,u)$ require one or more of the following to be bounded: $dP/dQ$, $dQ/dP$, $dP_{X}/dQ_{X}$, or $dQ_{X}/dP_{X}$. If $Q=P$ then these functions are all equal to unity, and so these boundedness conditions certainly hold. Given this, when composing these primitives to differentiate $\psi$ at $P$ we generally recommend choosing $Q=P$ to index the ambient Hilbert spaces.
In Appendix D.5, we describe how to transform primitives for conditional quantities to those for marginal quantities. This provides a way to derive a marginal mean primitive, for example. We also describe how to adapt the primitives from nonparametric to semiparametric models.
@@ 2.3 Example: nonparametric $R^{2}$
We express $\psi(P^{\prime}):=1-E_{P^{\prime}}[\{Y-\mu_{P^{\prime}}(X)\}^{2}]/\mathrm{Var}_{P^{\prime}}(Y)$ as a composition of eight primitives: a constant $\theta_{1}(P^{\prime},0)=(z\mapsto y)\in L^{2}(P)$, variance $\theta_{2}(P^{\prime},h_{1})=\mathrm{Var}_{P^{\prime}}[h_{1}(Z)]\in\mathbb{R}$, constant $\theta_{3}(P^{\prime},0)=(z\mapsto y)\in L^{2}(P)$, conditional mean $\theta_{4}(P^{\prime},h_{3})(\cdot)=E_{P^{\prime}}[h_{3}(Z)\,|\,X=\cdot\,]\in L^{2}(P_{X})$, lifting $\theta_{5}(P^{\prime},h_{4})=(z\mapsto h_{4}(x))\in L^{2}(P)$, pointwise operation $\theta_{6}(P^{\prime},h_{5})=(z\mapsto[y-h_{5}(z)]^{2})\in L^{2}(P)$, mean $\theta_{7}(P^{\prime},h_{6})=E_{P^{\prime}}[h_{6}(Z)]\in\mathbb{R}$, and differentiable function $\theta_{8}(P^{\prime},(h_{2},h_{7}))=1-h_{7}/h_{2}\in\mathbb{R}$. The redundancy of the constant maps $\theta_{1}$ and $\theta_{3}$ has no impact on the applicability of Algorithm 2. For the composition of primitives to express $\psi$, it must return $\psi(P^{\prime})$ for every $P^{\prime}\in\mathcal{M}$. This will hold if each claimed element of $L^{2}(P)$ or $L^{2}(P_{X})$ is a $P$-a.s. equivalence class of functions that satisfies an appropriate second-moment condition. To this end, we suppose all pairs of distributions in $\mathcal{M}$ are mutually absolutely continuous, which makes it so, for each $P^{\prime}\in\mathcal{M}$, $P^{\prime}$-a.s. and $P$-a.s. equivalence classes of functions are one and the same. For the moment condition, we suppose there exists $c<\infty$ such that $P^{\prime}(|Y|\leq c)=1$ for all $P^{\prime}\in\mathcal{M}$. To avoid dividing by zero, we also suppose $\mathrm{Var}_{P^{\prime}}(Y)>0$ for all $P^{\prime}$.
@@
[…]

### v3

@@ Abstract
Keywords: efficient influence function, asymptotic efficiency, pathwise differentiability, Hadamard differentiability, automatic differentiation, probabilistic programming, differentiable programming
@@ 1.1 Motivation
There is a long history of constructing estimators that are asymptotically efficient (‘efficient’ hereafter) in infinite-dimensional models (Levit, 1975; Hasminskii and Ibragimov, 1979; Pfanzagl, 1990; Bickel et al., 1993). Several approaches for doing this permit the use of arbitrary methods for estimating needed nuisance functions. These include one-step estimation (Pfanzagl, 1982), estimating equations (van der Laan and Robins, 2003), targeted learning (van der Laan and Rubin, 2006), and double machine learning (Chernozhukov et al., 2017). These frameworks construct efficient estimators using an object known as the efficient influence function (EIF). Though EIFs have been derived in many examples (Newey, 1990; Bickel et al., 1993; van der Laan and Robins, 2003), computing them in new problems can be challenging, requiring tools from functional analysis and differential geometry. These calculations can represent a bottleneck in scientific discovery, as some novel analyses cannot be performed until an expert derives the EIF.
@@
Automatic differentiation avoids these challenges (Wengert, 1964; Speelpenning, 1980). In this approach, a function $f$ is written as a composition of differentiable functions with known gradients, which we refer to as primitives. Then, an algorithm can iteratively apply the chain rule to derive the gradient of $f$ at a specified point. Unlike numerical differentiation, automatic differentiation is hyperparameter-free and is always accurate up to working precision. Differentiable programming extends automatic differentiation by identifying a compositional structure for $f$ directly via computer code that can be used to evaluate it, simplifying implementation (Baydin et al., 2018; Blondel and Roulet, 2024). Automatic differentiation and differentiable programming have been used with success in applications (Bischof, 1995; Homescu, 2011), and have played a crucial role in the rapid progress made in deep learning (Goodfellow et al., 2016). However, they have yet to be used to compute a semi- or nonparametric EIF. This may be due to the fact that EIFs correspond to a gradient of a functional defined on an infinite-dimensional statistical manifold, and so it has been unclear what primitives can be chained together to compute them algorithmically.
@@ 1.2 Contributions
    In Section 2.1, we provide an automatic differentiation algorithm for computing the EIF of any parameter that can be expressed as a composition of differentiable Hilbert-valued primitives.
@@
    In Section 2.2, we establish the differentiability of a variety of primitives. These primitives can be composed to express many parameters of interest.
We then turn to estimation, building on developments in probabilistic programming (van de Meent et al., 2018). Probabilistic programming and finite-dimensional automatic differentiation have been employed for Bayesian inference (Ge et al., 2018; Tran et al., 2018; Stan Development Team, 2023), parametric maximum likelihood estimation (Fournier et al., 2012; Kristensen et al., 2015), M-estimation (Zivich et al., 2022; Kosmidis and Lunardon, 2024), and efficient estimation in high-dimensional parametric models (Agrawal et al., 2024). However, to date there is not a probabilistic programming framework for efficiently estimating a generic smooth parameter that may rely on nonparametric nuisances like density and regression functions. Our remaining contributions address this gap.
@@
    In Section 3, we provide an algorithm to estimate any parameter that is expressed as a composition of known primitives and conditions under which this estimator will be efficient.
@@
    In Section 1.3, we introduce a proof-of-concept Python package implementing our approach, available at http://github.com/alexluedtke12/pydimple.
@@
The goal of dimple is similar in spirit to that of automatic debiased machine learning methods: to avoid analytical derivations when constructing efficient estimators (Chernozhukov et al., 2022a, b, c). Those existing methods apply to regression functionals, whose EIFs depend on the functional’s form through nuisance functions known as Riesz representers. Automatic debiased machine learning provides loss functions that can be used to estimate these nuisances without explicitly deriving their form. Compared to these methods, dimple innovates in several directions. It allows the functional to depend on the distribution through non-regression summaries, uses automatic differentiation to compute the EIF, and is integrated into a user-friendly probabilistic programming framework. It also applies to general Hilbert-valued parameters (Luedtke and Chung, 2024).
Beyond simplifying estimation, our automatic differentiation approach and the chain rule that underlies it may be of independent interest for their use in deriving semiparametric efficiency bounds, which are given by the variance of the EIF when it exists (van der Vaart and Wellner, 1989; van der Vaart, 1991a). Our theoretical results show this EIF exists for any parameter expressible as a composition of differentiable primitives, and applying automatic differentiation by hand gives its analytic form. An example of this is given in Section 2.3.
@@ 1.3 Illustrations of proposed approach
Our proof-of-concept Python package, pydimple, uses dimple to construct nonparametric efficient estimators. We illustrate it with three examples. The following Python code imports the classes and functions from the package needed to run them:
from pydimple import Distribution, E, Density, Var, RV, estimate
@@ 1 P = Distribution(data=dat)
The first line of code defines a distribution $P$ and identifies a dataset dat containing iid draws from it to be used for estimation. The next two lines serve as a blueprint of the statistical estimation problem, formulating it using primitives from pydimple. The fourth line computes a 5-fold cross-fitted estimator, Wald-type standard error, and 95% confidence interval of $\psi(P)$. Internally, pydimple estimates density functions using kernel density estimation and all other nuisances using lightgbm (Ke et al., 2017), with hyperparameters tuned via Optuna (Akiba et al., 2019).
@@ 4 R2 = 1-E(P,(RV(’Y’)-mu)**2)/v # $\psi(P)$
The final example concerns the estimation of the longitudinal G-formula (Robins, 1986; Bang and Robins, 2005; Luedtke et al., 2017; Rotnitzky et al., 2017; Chernozhukov et al., 2022b). This parameter is defined in longitudinal settings where covariate-treatment pairs $(X_{t},A_{t})$ are collected at times $t=0,1,\ldots,T-1$, and then an outcome $Y$ is collected at the end of the study. For $t\in\{0,1,\ldots,T-1\}$, let $H_{t}$ denote the history $(X_{0},A_{0},X_{1},A_{1},\ldots,X_{t-1},A_{t-1},X_{t})$ available just before treatment $A_{t}$ is administered. Also let $H_{T}=(H_{T-1},A_{T},Y)$ denote all of the covariate, treatment, and outcome data. The parameter $\psi(P)$ is defined recursively by letting $\mu_{P,T}(h_{T})=y$, $\mu_{P,t}(h_{t})=E_{P}[\mu_{P,t+1}(H_{t+1})\,|\,A_{t}=1,H_{t}=h_{t}]$ for $t=T-1,T-2,\ldots,0$, and $\psi(P)=E_{P}[\mu_{P,0}(H_{0})]$. Under causal conditions, $\psi(P)$ gives the counterfactual mean outcome when everyone receives treatment $1$ at all time points. The recursive definition of this parameter can be easily expressed in pydimple via a for loop, as illustrated below.
@@ 5 H = [f’X{j}’ for j in range(t+1)]+[f’A{j}’ for j in range(t)] # $H_{t}$
Example 3: Longitudinal G-Formula (Robins, 1986).
In Section 4, we evaluate the performance of pydimple in a Monte Carlo study. The results support that dimple can produce asymptotically efficient estimators and valid confidence intervals. At sample sizes of $n\geq 1000$, the confidence intervals have approximately nominal coverage and nearly optimal widths, scaling with the standard deviation of the EIF. These results are not surprising given that dimple uses a variant of one-step estimation, which has been shown to perform well in a variety of settings (Pfanzagl, 1982; Bickel et al., 1993). Details on this estimation framework and its theoretical properties are given in Section 3.1.
@@ 2.1 Algorithm and theoretical guarantee
Though we have not seen this precise notion of differentiability introduced previously, total pathwise differentiability is closely related to classical notions that apply to functions of a single argument, $u\in\mathcal{U}$ or $P\in\mathcal{M}$. In particular, the total pathwise differentiability of $\theta$ at $(P,u)$ implies $\theta(P,\,\cdot\,)$ is Hadamard differentiable (Averbukh and Smolyanov, 1967) and $\theta(\,\cdot\,,u)$ is pathwise differentiable (van der Vaart, 1991a), and the converse implication holds under additional regularity conditions — see Lemmas S5 and S6 in the appendix. In the special case where $\mathcal{U}$ is the trivial vector space $\{0\}$, $\theta$ is totally pathwise differentiable if and only if $\nu(\cdot):=\theta(\,\cdot\,,0)$ is pathwise differentiable. In these cases, $\dot{\theta}_{P}^{*}(w)=(\dot{\nu}_{P}^{*}(w),0)$. The map $\dot{\nu}_{P}^{*}:\mathcal{W}\rightarrow\dot{\mathcal{M}}_{P}$ is called the efficient influence operator of $\nu$ at $P$; it can be used to draw inferences about $\nu(P)$ (Luedtke and Chung, 2024). When $\mathcal{W}=\mathbb{R}$ and $w=1$, $\dot{\nu}_{P}^{*}(w)$ is known as the EIF of $\nu$ at $P$ (Bickel et al., 1993, Definition 5.2.3).
Like Hadamard and pathwise differentiability (Averbukh and Smolyanov, 1967; van der Vaart, 1991b; Kennedy, 2022), total pathwise differentiability satisfies a chain rule — see Lemma S3 in the appendix. The availability of a chain rule suggests a form of automatic differentiation (Linnainmaa, 1970), which we present in Algorithm 2. Initially, this algorithm requires the evaluation of the composition that defines $\psi$. Subsequently, information backpropagates through the indices $j$ of the composition to evaluate the efficient influence operator. This information is derived from the adjoint $\dot{\theta}_{j,P,h_{\mathrm{pa}(j)}}^{*}$ of the differential operator of $\theta_{j}$ at $(P,h_{\mathrm{pa}(j)})$, where $h_{\mathrm{pa}(j)}$ is as defined as in Algorithm 1 when that algorithm is called with input $P$.
@@ 2.2 Useful primitives
Table 1 provides a selection of useful primitives. In Appendix C, we establish that, under conditions, each is totally pathwise differentiable when $\mathcal{M}$ is a locally nonparametric model of distributions on $\mathcal{Z}$. Some of these primitives involve a coarsened random variable $X=\mathscr{C}(Z)$, where $\mathscr{C}:\mathcal{Z}\rightarrow\mathcal{X}$ is a many-to-one map; when $Z\sim Q\in\mathcal{M}$, we let $Q_{X}$ denote the marginal distribution of $X$. The primitives fall into three distinct categories, which we elaborate on below.
Table 1: Primitives $\theta$ with domain ×$\mathcal{M}\times\mathcal{U}$. Given the specified ambient Hilbert spaces $\mathcal{T}\supseteq\mathcal{U}$ and $\mathcal{W}\supseteq\mathrm{Image}(\theta)$, each is totally pathwise differentiable under conditions given in Appendix C.
@@
The choice of ambient Hilbert spaces $\mathcal{T}$ and $\mathcal{W}$ can impact a primitive’s total pathwise differentiability. For the first five primitives in Table 1 and the conditional density primitive, this choice involves selecting a $Q\in\mathcal{M}$ indexing one or more $L^{2}$ spaces. As detailed in Appendix C, sufficient conditions for each primitive’s total pathwise differentiability at a given $(P,u)$ require one or more of the following to be bounded: $dP/dQ$, $dQ/dP$, $dP_{X}/dQ_{X}$, or $dQ_{X}/dP_{X}$. If $Q=P$ then these functions are all equal to unity, and so these boundedness conditions certainly hold. Given this, when composing these primitives to differentiate $\psi$ at $P$ we generally recommend choosing $Q=P$ to index the ambient Hilbert spaces.
In Appendix C.5, we describe how to transform primitives for conditional quantities to those for marginal quantities. This provides a way to derive a marginal mean primitive, for example. We also propose a way to adapt the primitives from nonparametric to semiparametric models, though we caution that this proposal requires projecting onto the tangent space of the model and so implementing it can require specialized expertise.
@@ 2.3 Example: nonparametric $R^{2}$
We express $\psi(P^{\prime}):=1-E_{P^{\prime}}[\{Y-\mu_{P^{\prime}}(X)\}^{2}]/\mathrm{Var}_{P^{\prime}}(Y)$ as a composition of eight primitives: a constant $\theta_{1}(P^{\prime},0)=(z\mapsto y)\in L^{2}(P)$, variance $\theta_{2}(P^{\prime},h_{1})=\mathrm{Var}_{P^{\prime}}[h_{1}(Z)]\in\mathbb{R}$, constant $\theta_{3}(P^{\prime},0)=(z\mapsto y)\in L^{2}(P)$, conditional mean $\theta_{4}(P^{\prime},h_{3})(\cdot)=E_{P^{\prime}}[h_{3}(Z)\,|\,X=\cdot\,]\in L^{2}(P_{X})$, lifting $\theta_{5}(P^{\prime},h_{4})=(z\mapsto h_{4}(x))\in L^{2}(P)$, pointwise operation $\theta_{6}(P^{\prime},h_{5})=(z\mapsto[y-h_{5}(z)]^{2})\in L^{2}(P)$, mean $\theta_{7}(P^{\prime},h_{6})=E_{P^{\prime}}[h_{6}(Z)]\in\mathbb{R}$, and differentiable function $\theta_{8}(P^{\prime},(h_{2},h_{7}))=1-h_{7}/h_{2}\in\mathbb{R}$. Here, $\mathrm{pa}(1)=\mathrm{pa}(3)=\emptyset$, $\mathrm{pa}(2)=\{1\}$, $\mathrm{pa}(4)=\{3\}$, $\mathrm{pa}(5)=\{4\}$, $\mathrm{pa}(6)=\{5\}$, $\mathrm{pa}(7)=\{6\}$, and $\mathrm{pa}(8)=\{2,7\}$. The redundancy of the constant maps $\theta_{1}$ and $\theta_{3}$ has no impact on the applicability of Algorithm 2. For the composition of primitives to express $\psi$, it must return $\psi(P^{\prime})$ for every $P^{\prime}\in\mathcal{M}$. This will hold if each claimed element of $L^{2}(P)$ or $L^{2}(P_{X})$ is a $P$-a.s. equivalence class of functions that satisfies an appropriate second-moment condition. To this end, we suppose all pairs of distributions in $\mathcal{M}$ are mutually absolutely continuous, which makes it so, for each $P^{\prime}\in\mathcal{M}$, $P^{\prime}$-a.s. and $P$-a.s. equivalence classes of functions are one and the same. For the moment condition, we suppose there exists $c<\infty$ such that $P^{\prime}(|Y|\leq c)=1$ for all $P^{\prime}\in\mathcal{M}$. To avoid dividing by zero, we also suppose $\mathrm{Var}_{P^{\prime}}(Y)>0$ for all $P^{\prime}$.